Pipe A can fill a tank in 4 hours, and pipe B can fill the tank in 3 hours. With the tank empty, pipe A is turned on. One hour later, pipe B is turned on. How long will pipe B run before the tank is filled?
Please help me with this questions, thanks.

Answer by jorel1380(3719)

If pipe A can fill a tank in 4 hours, and it is left on for an hour, it will have filled 1/4 of the tank. Together , the pipes can fill 1/4+1/3 tank an hour.
----------------
(1/4)t+(1/3)t=3/4
Multiply both sides by 12
3t+4t=9
7t=9
t=9/7 hours
Pipe B will run along with pipe A together for 9/7 hours before they fill the tank.